OpenStudy (anonymous):

11.Graph the following equations and state what the solution is? Y = 2x + 4 y = -x – 1

OpenStudy (triciaal):

EQN 1 WHEN X = 0 Y = 4 when y = 0 x = -2 eqn 2: when x = 0 y = -1 when y =0 x= -1 when they meet 2x + 4 = -x -1 solve for x substitute in either eqn to get y

OpenStudy (anonymous):

3x = -5

OpenStudy (triciaal):

so x = -5/3 now find y

OpenStudy (triciaal):

always check your solution as the final step

OpenStudy (anonymous):

how do I find Y

OpenStudy (triciaal):

from original eqn you have x

OpenStudy (anonymous):

so would that be 2x +4 = -5/3 -1

OpenStudy (nikato):

you could also substitute -5/3 for x in one of ur original equation

OpenStudy (triciaal):

y = 2x + 4 or y = -x -1

OpenStudy (nikato):

What u did will not help u because you have already found x

OpenStudy (anonymous):

y = -5/3x - 1

OpenStudy (nikato):

It's actually y=-(-5/3)-1 So y=5/3 -1

OpenStudy (triciaal):

y=2(-5/3) + 4 -10/3 +4 =(-10 +12)/3 =2/3

OpenStudy (anonymous):

is the answer 2/3

OpenStudy (jdoe0001):

\(\bf {\color{red}{ Graph}}\) the following equations and state what the solution is?

OpenStudy (anonymous):

is 2/3 the slope

OpenStudy (triciaal):

you should notice it doesn't matter which one you substitute in get the same result

OpenStudy (jdoe0001):

pick a couple of random "x"'s for each to get a "y", draw them with the 2 points found

OpenStudy (anonymous):

what the X value and whats the Y value

OpenStudy (triciaal):

2/3 is not the slope we were solving for y

OpenStudy (anonymous):

is y the slope

OpenStudy (anonymous):

no sorry I meant os 2/3 = y

OpenStudy (triciaal):

did you draw the graph from the points obtained when x = 0 and when y = 0?

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

like jdoe0001 said above you could use random values of x find the y to get points for the graph.

OpenStudy (anonymous):

like X could be 4

OpenStudy (whpalmer4):

Easy way to graph straight lines in a problem like this is to find the x and y intercepts by substituting 0 for one and solving for the other. Put a ruler down on those two points and draw a straight line. I believe that is what @triciaal was doing in her first two lines of her first post. No need to find any other points to graph the straight lines — two points will determine the line and plugging in 0 is likely to be as easy as anything.

OpenStudy (whpalmer4):

So for the first equation: \[y = 2x + 4\] Plug in \(x=0\): \[y = 2(0)+4\]\[y = 4\]First point is \((0,4)\) (that's the y-intercept, the point at which the line crosses the y-axis) Plug in \(y = 0\): \[0 = 2x+4\]\[-2x=4\]\[x=-2\]Second point is \((-2,0)\) (the x-intercept)